OpenStudy (anonymous):

x^4 *y^4 = 16, normal (2,1). Find the slope of the the line that is normal to the curve. A) y = 1/8x B) y = -1/2 x +2 C) y = -2x+5 D) y = 2x-3

OpenStudy (binary3i):

y'=Y/X=1/2 this is the slope of tangent at the point (2,1) slope of the normal will be -(1/y') so slope of normal is -2 equation becomes Y=-2X + c ut the values of X and Y 1=-4+c c=5 therefore equation is Y=-2X+5